OpenStudy (anonymous):

what is 11/8 as a decimal?? please help

OpenStudy (timmyk14):

1.375

OpenStudy (alexandervonhumboldt2):

1000/8=125 11*125 1.375 no direct answers

OpenStudy (cwrw238):

@Timmyk14 - plz don't just give the answers . The moderators wont be pleased. Check out Code of Conduct

OpenStudy (anonymous):

ok but for this problem says...what is the volume to the nearest tenth of an inch..for a cone that has a diameter of 2 3/4 and a height of 4

OpenStudy (anonymous):

the anser is 7.9

OpenStudy (anonymous):

and I keep getting 7.123875

OpenStudy (alexandervonhumboldt2):

1.375

OpenStudy (anonymous):

ok.but what would you do on that problem?? would you convert the fractions to decimals or keep doing the problems as a fraction

OpenStudy (sleepyjess):

Volume of a cone is \(\pi r^2\dfrac{h}{3}\). The height (h) is given as 4 and the diameter is given as 2.75, but we need to divide this by 2 to get the radius.

OpenStudy (sleepyjess):

Does that make sense so far @lopez09 ?

OpenStudy (sleepyjess):

I prefer to convert the fraction to a decimal